OpenStudy (anonymous):

Graph the equation y = x2 - 3x -4. For what value or values of x is y = 0

OpenStudy (anonymous):

This is asking for the solutions to the quadratic because the solutions are where y = 0. Set y = 0. \[0 = x^2 - 3x - 4\]Can you factor this?

OpenStudy (anonymous):

x2 messing me up is it just x+x-3x-4

OpenStudy (anonymous):

No. \[x^2 = x \times x\]

OpenStudy (anonymous):

What 2 factors of -4 add up to -3?

OpenStudy (anonymous):

-4+1

OpenStudy (anonymous):

Alright. Place the factors into their positions. \[x^2 - 3x - 4 = 0\implies (x - 4)(x + 1) = 0\]Now, remember how 0 times anything equals 0?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Alright. That's one value. x = 4. Now do the same for the other part. x + 1 = 0 x = ?

OpenStudy (anonymous):

sorry this is times so I'm super stressed

OpenStudy (anonymous):

No problem. x = 4, -1 For those values, y is equal to 0.

OpenStudy (anonymous):

thank you so much for walking me through it

OpenStudy (anonymous):

np :)